OpenStudy (anonymous):

Did i get these right.. What is the value of the y variable in the solution to the following system of equations? 3x + y = 14 -3x + 2y = 1 i got 5 What is the value of the x variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 i got -2 What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 i got no solution

OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations eliminated today? i got 412

OpenStudy (anonymous):

3x + y = 14 -3x + 2y = 1 X = 3 Y = 5 @Courage. Good :)

OpenStudy (anonymous):

k good so far

OpenStudy (anonymous):

3x − 6y = 3 7x − 5y = −11 X = -3 Y = -2 @Courage. Good :)

OpenStudy (anonymous):

2x − y = 6 4x − 2y = 12 INFINITE SOLUTIONS @Courage. Not Correct :(

OpenStudy (anonymous):

oh ok